OpenStudy (anonymous):

How do you graph xy=15? I know that 15 is k or the inverse variation but i still keep messing up my x and y points...

OpenStudy (anonymous):

Would it help if I retyped xy=15 to y=15/x?

OpenStudy (anonymous):

no, i did that a couple of times and i dont think i got the right answers. What to you do after that?

OpenStudy (anonymous):

you just start plugging in values of x and getting the y value and graph those points

OpenStudy (anonymous):

Do you start with any number?

OpenStudy (anonymous):

whichever number you want to, it's just like graphing regular lines. Just to let you know if you're doing it right, your graph should look something like this depending on the interval on which you're graphing: 

OpenStudy (phi):

I would write it y = 15/x now pick some x values. at x=1 y= 15/1 = 15 at x= 15 y= 15/15 = 1 at x= 30 y = 15/30 = 1/2 at x= 1/2 y= 15/(1/2)= 30 so plot the points (1/2,30) (1,15) (15,1) (30, 1/2) after you plot the points connect the dots now try the same x numbers, but make them negative, and find the y numbers.